# Hysteresis Compensation of Flexible Continuum Manipulator using RGBD Sensing and Temporal Convolutional Network

### v2

@@ Abstract
Flexible continuum manipulators are valued for minimally invasive surgery, offering access to confined spaces through nonlinear paths. However, cable-driven manipulators face control difficulties due to hysteresis from cabling effects such as friction, elongation, and coupling. These effects are difficult to model due to nonlinearity and the difficulties become even more evident when dealing with long and coupled, multi-segmented manipulator. This paper proposes a data-driven approach based on Deep Neural Networks (DNN) to capture these nonlinear and previous states-dependent characteristics of cable actuation. We collect physical joint configurations according to command joint configurations using RGBD sensing and 7 fiducial markers to model the hysteresis of the proposed manipulator. Result on a study comparing the estimation performance of four DNN models show that the Temporal Convolution Network (TCN) demonstrates the highest predictive capability. Leveraging trained TCNs, we build a control algorithm to compensate for hysteresis. Tracking tests in task space using unseen trajectories show that the proposed control algorithm reduces the average position and orientation error by 61.39% (from $\mathbf{13.7mm}$ to $\mathbf{5.29mm}$) and 64.04% (from ${}^{\circ}$ to ${}^{\circ}$), respectively. This result implies that the proposed calibrated controller effectively reaches the desired configurations by estimating the hysteresis of the manipulator. Applying this method in real surgical scenarios has the potential to enhance control precision and improve surgical performance.
@@ I Introduction
Height between notches, h
$\mathrm{0.5mm}$
Thickness of the disk, l
$\mathrm{0.7mm}$
Round radius of the hinge, R
$\mathrm{0.35mm}$
Front thickness, t
$\mathrm{0.4mm}$
Side thickness of the hinge, b
$\mathrm{1.3mm}$
Radius of module, r
$\mathrm{2.4mm}$
@@ III-A Design Considerations
Flexure hinge-based continuum manipulators offer the advantage of miniaturization due to their simple structure.The specific design parameters are detailed in Table I. (See Fig. 2). Finite element analysis was conducted to confirm the suitability of flexure hinge design made of PEEK. The bending angle of a single hinge model is ${}^{\circ}$, with a maximum strain of 0.0378. Additionally, when the deformation amplitude ranges from 0.04 to 0.035 at $\mathrm{0.5Hz}$ between 0.04 and 0.035, the number of repetitions ranges from 18454 to 92078 [27]. This generally corresponds to a high cycle life, ranging from $10^{4}$ to $10^{6}$. The proposed manipulator consists of a total of 11 hinge modules, maintaining a high cycle life even with a ${}^{\circ}$ deflection angle.
@@ IV-B Dataset Collection and Hysteresis Observation
The $\textbf{q}_{\mathrm{cmd}}$ was generated through a two-step process. We first randomly selected 1,802 joint configurations within the manipulator’s joint limits. These configurations represent a diverse set of poses within reachable workspace. Next, to ensure a uniform distance ($3\sqrt{5}$) of data points across the joint space, interpolation was performed between these randomly chosen configurations, and resulting 30,707 command joint angles. The interpolation method utilized through (16).
$\textbf{q}_{\mathrm{p}_{i,j}}=\textbf{q}_{\mathrm{p}_{i}}+j\cdot\frac{\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}}{\frac{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}{3\sqrt{5}}}\>\>\>\>(j\leq\lfloor\frac{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}{3\sqrt{5}}\rfloor)$ (32)
where $\textbf{q}_{\mathrm{p}_{i}}$ and $\textbf{q}_{\mathrm{p}_{i+1}}$ are consecutive randomly chosen configuration, $\lfloor\frac{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}{3\sqrt{5}}\rfloor$ is the number of interpolated points between them, |||||$\cdot$||| | ⋅ | | is the $l_{2}$ norm.
@@
  Joint angle
$q_{1}$[${}^{\circ}$] $q_{2}$[${}^{\circ}$] $q_{3}$[${}^{\circ}$] $q_{4}$[${}^{\circ}$] $q_{5}$[${}^{\circ}$]
MAE 16.31 10.09 11.21 12.02 13.84
SD 11.59 7.67 7.56 8.91 10.37
@@ IV-C Hysteresis Modeling using Deep Learning
[…]
@@
[…]
@@ VI Conclusions
$q_{1}$[${}^{\circ}$]
$q_{2}$[${}^{\circ}$]
$q_{3}$[${}^{\circ}$]
$q_{4}$[${}^{\circ}$]
$q_{5}$[${}^{\circ}$]
Uncalibrated control MSE
15.60
13.73
12.97
11.37
18.97
SD
11.42
10.54
8.50
8.19
12.15

### v3

@@ Abstract
Flexible continuum manipulators are valued for minimally invasive surgery, offering access to confined spaces through nonlinear paths. However, cable-driven manipulators face control difficulties due to hysteresis from cabling effects such as friction, elongation, and coupling. These effects are difficult to model due to nonlinearity and the difficulties become even more evident when dealing with long and coupled, multi-segmented manipulator. This paper proposes a data-driven approach based on Deep Neural Networks (DNN) to capture these nonlinear and previous states-dependent characteristics of cable actuation. We collect physical joint configurations according to command joint configurations using RGBD sensing and 7 fiducial markers to model the hysteresis of the proposed manipulator. Result on a study comparing the estimation performance of four DNN models show that the Temporal Convolution Network (TCN) demonstrates the highest predictive capability. Leveraging trained TCNs, we build a control algorithm to compensate for hysteresis. Tracking tests in task space using unseen trajectories show that the proposed control algorithm reduces the average position and orientation error by 61.39% (from $\mathbf{13.7mm}$ to $\mathbf{5.29mm}$) and 64.04% (from 31.17 to 11.21), respectively. This result implies that the proposed calibrated controller effectively reaches the desired configurations by estimating the hysteresis of the manipulator. Applying this method in real surgical scenarios has the potential to enhance control precision and improve surgical performance.
Index Terms:
Tendon/Wire Mechanism, Machine Learning for Robot Control, Modeling, Control, and Learning for Soft Robots
@@ I Introduction
                 Height between notches, h               $\mathrm{0.5mm}$
                 Thickness of the disk, l               $\mathrm{0.7mm}$
                 Round radius of the hinge, R               $\mathrm{0.35mm}$
                 Front thickness, t               $\mathrm{0.4mm}$
                 Side thickness of the hinge, b               $\mathrm{1.3mm}$
                 Radius of module, r               $\mathrm{2.4mm}$
@@ III-A Design Considerations
Flexure hinge-based continuum manipulators offer the advantage of miniaturization due to their simple structure.The specific design parameters are detailed in Table I. (See Fig. 2). Finite element analysis was conducted to confirm the suitability of flexure hinge design made of PEEK. The bending angle of a single hinge model is 5.5, with a maximum strain of 0.0378. Additionally, when the deformation amplitude ranges from 0.04 to 0.035 at $\mathrm{0.5Hz}$ between 0.04 and 0.035, the number of repetitions ranges from 18454 to 92078 [25]. This generally corresponds to a high cycle life, ranging from $10^{4}$ to $10^{6}$. The proposed manipulator consists of a total of 11 hinge modules, maintaining a high cycle life even with a 60.5 deflection angle.
@@ IV-B Dataset Collection and Hysteresis Observation
The $\textbf{q}_{\mathrm{cmd}}$ was generated through a two-step process. We first randomly selected 1,802 joint configurations within the manipulator’s joint limits. These configurations represent a diverse set of poses within reachable workspace. Next, to achieve a uniform distribution with an interval of $3^{\circ}$ on each of the 5 joints across joint space (resulting in an euclidean distance of $3\sqrt{5}$ between points), interpolation was performed between these randomly chosen configurations, and resulting 30,707 command joint angles. The interpolation method utilized through (16).
$\textbf{q}_{\mathrm{p}_{i,j}}=\textbf{q}_{\mathrm{p}_{i}}+3j\sqrt{5}\cdot\frac{\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}}{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}\>\>\>\>(j\leq\lfloor\frac{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}{3\sqrt{5}}\rfloor)$ (16)
where $\textbf{q}_{\mathrm{p}_{i}}$ and $\textbf{q}_{\mathrm{p}_{i+1}}$ are consecutive randomly chosen configuration, $\lfloor\frac{||\textbf{q}_{\mathrm{p}_{i+1}}-\textbf{q}_{\mathrm{p}_{i}}||}{3\sqrt{5}}\rfloor$ is the number of interpolated points between them, and |||||$\cdot$||| | ⋅ | | is the $l_{2}$ norm.
@@ IV-C Hysteresis Modeling using Deep Learning
1) Approaches: we employ the forward approach $(\textit{f}_{\mathrm{\theta}})$ and the inverse approach $({\textit{f}}_{\theta}^{\>\>-1})$. The forward approach predicts physical joint angle from sequence of command joint angles while the inverse approach predicts command joint angle from sequence of physical joint angles. In mathematical terms,
@@
2) Architectures: We consider the following candidate architectures for $\textit{f}_{\mathrm{\theta}}$ and $\textit{f}_{\mathrm{\theta}}^{\>-1}$.
@@ VI Conclusions
$q_{1}$[] $q_{2}$[] $q_{3}$[] $q_{4}$[] $q_{5}$[]
Uncalibrated control MSE 15.60 13.73 12.97 11.37 18.97
SD 11.42 10.54 8.50 8.19 12.15